# Robust Utility Optimization via a GAN Approach

### v1

@@ 4 Experiments
We start with experiments in easy settings, where an analytic solution exists in Section 4.1. Then, in Sections 4.2 and 4.3, we introduce friction, different utility functions and more realistic path-dependent penalization. In all of these cases, analytic solutions are not known. Therefore, in Section 4.3.3, we introduce different metrics to evaluate the performance of our model. Moreover, in Section 4.4, we leverage our GAN-based algorithm in an experimental study of small, proportional trading costs on optimal investments in robust and non-robust utility optimization.
@@
4.1 Friction-less Settings with Analytic Solutions
In the following examples we always use stock markets with risk free interest rate $r=1.5\%$, step size $\delta_{t}=\frac{1}{65}$ and $65$ steps. All stocks start at the initial value $S_{0}=1$ and are given by an Itô-diffusion driven by one or multiple Brownian motions. We sample a total of $200$K paths of Brownian motions, which are used for all experiments in this section such that there is no random influence due to different samples across different algorithm evaluations, with an $80/20$ train/test split.
We use the analytic solution for the market volatility $\Sigma^{*}$ (and $\mu^{*}$ if the drift is also robustified) to get an evaluation metric that computes the expected utility of either the learned or the analytic investment strategy against the analytic solution for the market. Let us define
@@ 4.1 Friction-less Settings with Analytic Solutions
then we compute $E_{u}(\pi,\Sigma^{*},\mu^{*})$ for the learned strategy $\pi=\pi^{\text{NN}}$ and for the analytic strategy $\pi=\pi^{*}$ (where $\mu^{*}=\mu$ in case the drift is fixed) and compare them via their relative error
@@ 4.1.1 Stock Market with 1 Risky Asset & Robust Volatility
[…]
where $\tilde{\sigma}\in\mathbb{R}_{+}$ is a constant reference value for the volatility (this could for example be estimated from historical data), and $\lambda_{1}$ is the scaling factor for the penalization. In this case there exists an analytic solution for the optimal actor and market strategy, which was given in Guo et al. (2022, Section 7.1). In particular, the optimal investment strategy and volatility are both constant and given as the solution of
@@
which we solve numerically to get a reference value. We use this setting to test and compare different models, since it offers an easy evaluation metric. In particular, we compute the expected utility (without penalty) when using the actor’s trained investment strategy and the analytic solution for the volatility (evaluation metric) and compare it to the expected utility when using the analytic solution for the actor and the market (reference evaluation metric). The reason for this type of evaluation procedure is that our ultimate goal is to achieve a good robust investment strategy from the GAN training, which should perform well in the harshest possible market environment.
@@ 4.1.2 Stock Market with Multiple Risky Assets & Robust Volatility
[…]
where $\Sigma_{t}=\sigma_{t}\sigma_{t}^{\top}$. The utility function is $u(x)=\ln(x)$. To derive the analytic solution of the (continuous-time) optimization problem (4), we first use Itô’s lemma to derive the continuous dynamics of $\ln(X_{T}^{\pi,\mu,\sigma})$. After dropping the martingale terms, (4) simplifies to
@@
We follow (Guo et al., 2022, Section 7.2) and compute the instantaneous derivatives of the integrand with respect to $\pi_{t}$ and $\Sigma_{t}$ and set them to 00. This leads (after some rearrangement) to the equation systems
@@
The analytic solutions $\pi^{*},\Sigma^{*}$ can then be computed by first substituting $\Sigma$ in the second equation and numerically solving it for $\pi$ to obtain $\pi^{*}$ and then using this to compute $\Sigma^{*}$.
@@
      negatively correlated, asymmetric setting: assets’ volatilities are $0.15$ and $0.35$, their correlation is $-0.9$, leading to the reference volatility
@@
In all cases, our model achieves a very small relative error compared to the analytic strategy $\pi^{*}$ satisfying $\operatorname{err}_{rel}^{\Sigma^{\star},\mu}(\pi^{\text{NN}},\pi^{\star})<0.2\%$.
@@ 4.1.3 Robustifying Volatility and Drift
In this section, we robustify against uncertainty in the drift in addition to uncertainty in the volatility. Analogously to the robustification w.r.t. the volatility of Section 4.1.2, we choose a reference value $\tilde{\mu}$ for the drift and consider the joint penalty function (5). Then, for log-utility $u=\ln$, we can compute the analytic solution as in Section 4.1.2. In particular, we first apply Itô’s lemma to (4), then drop the martingale terms and compute the instantaneous derivative of the integrand with respect to $\pi_{t},\Sigma_{t}$ and $\mu_{t}$. Setting them to 00 yields
$\Sigma=\frac{\pi\pi^{\top}}{4\lambda_{1}}+\Sigma^{0}\quad\text{and}\quad\Sigma\pi=(\mu-r)\quad\text{and}\quad\pi=2\lambda_{2}(\mu^{0}-\mu),$
which can be solved numerically for the unique analytic solution $\pi^{*},\Sigma^{*},\mu^{*}$. Therefore, we can again compute the relative errors $\operatorname{err}_{rel}^{\Sigma^{\star},\mu^{*}}(\pi^{\text{NN}},\pi^{\star})$.
@@
and use balanced penalty weights $\lambda_{1}=\lambda_{2}=1$ (cf. Equation 5). In all combinations, we find a relative error $\operatorname{err}_{rel}^{\Sigma^{\star},\mu^{*}}(\pi^{\text{NN}},\pi^{\star})<0.2\%$.
Moreover, we get similar results when replacing the recurrent neural network (RNN) structures of generator and discriminator by standard feed-forward neural networks (FFNN) (that are reused for each time step). This was anticipated, since the optimal analytic strategy is constant and therefore not path dependent. In Figure 1, we show sample paths of the trained trading strategy evaluated against the analytic solution $\Sigma^{\star},\mu^{\star}$ for the market, and against the trained discriminator for the market on the test dataset. We see that in both cases, the trained strategy resembles the analytic strategy quite well. Interestingly, we further observe that the trained discriminator produces volatilities fairly similar to the analytic solution $\Sigma^{\star}$ (bottom left panels in Figure 1. This is not the case for adversarially learned drift values that differ from $\mu^{\star}$ (top right panels in Figure 1).
Figure 1: Test sample paths of the trained trading strategy evaluated against the analytic solution for the market (left) and against the trained discriminator for the market (right). A FFNN is used for generator and discriminator.
@@ 4.2 Settings with Proportional Transaction Costs
Since analytic solutions $\pi_{c}^{*},\Sigma_{c}^{*}$ (and $\mu^{*}$) for the problem including transaction costs are not known, we instead use the analytic solution for the corresponding problem without transaction costs (problem (15)), which can be considered as the best strategy that can be derived analytically. In contrast to this limitation of closed-form solutions, a simple adjustment of the training framework allows us to train our GAN model in the market with transaction costs. As before, we compute the relative errors $\operatorname{err}_{rel}^{\Sigma^{\star},\mu^{*}}(\pi^{\text{NN}},\pi^{\star})$, where market friction is in place. Clearly, the analytic market parameters are, in general, not the optimizers of the problem with friction, however, since our goal is to achieve a robust strategy, it is still reasonable to test the strategies against these (maybe suboptimal) market parameters. We note again that negative values correspond to superior performance of the GAN-based method.
We consider the same settings for the reference volatility $\tilde{\Sigma}$ and reference drift $\tilde{\mu}$ as in Section 4.1.3. For small transaction costs $c_{\text{prop}}=0.1\%$, we get $\operatorname{err}_{rel}^{\Sigma^{\star},\mu^{*}}(\pi^{\text{NN}},\pi^{\star})\in(-2.3\%,-0.3\%)$, which indicates slight outperformance of the (friction-aware) GAN model compared to the analytic solution $\pi^{\star}$ of the friction-less case. In particular, we see that taking friction into account improves the performance of the strategy, but that the difference is not very large due to the small size of the transaction costs. For larger transaction costs $c_{\text{prop}}=1\%$, this effect becomes considerably stronger, with relative errors resulting in $\operatorname{err}_{rel}^{\Sigma^{\star},\mu^{*}}(\pi^{\text{NN}},\pi^{\star})\in(-165.3\%,-10.5\%)$.
@@ 4.3 Realistic Settings for Generic Penalized Robust Utility Optimization
In this section, we analyze what we consider the most realistic setting for penalized robust utility optimization: We robustify against uncertainty in both the volatility and the drift, in a market with transaction costs, where we consider general power utilities corresponding to various preference structures including but not limited to log-preferences. Moreover, we study the path-wise penalties $R_{1}$ and $R_{2}$ of Equations 7 and 6. These penalties are more realistic than the penalties of Sections 4.1 and 4.2, since $R_{1}$ and $R_{2}$ can be estimated well from observed stock paths in contrast to the (empirically) unobservable instantaneous volatility and drift parameters that appear in Equation 14.
In this general setting, no analytic solution to the robust utility optimization problem is known. Moreover, this setting differs in too many regards from the setting in Section 4.1, where only log utility, no transaction costs and instantaneous penalization were considered, such that it is not reasonable to benchmark against the analytic solutions of Section 4.1 anymore. Therefore, we introduce new metrics to evaluate model performance in Section 4.3.3. Prior to that, we define general power utility functions and realistic path-wise penalties in Sections 4.3.1 and 4.3.2.
@@ 4.3.1 Power Utility as Generalization of Log Utility
where $p\in\mathbb{R}_{\geq 0}\setminus\{1\}$. In the limit $p\to 1$ the $\log$-utility function is recovered, hence we write $u_{1}:=\ln$, and for $p=0$ $u_{p}$ equals the risk-neutral utility $id-1$. In particular, $p\in(0,1)$ leads to less risk aversion compared to the $\log$-utility function, while $p>1$ leads to more risk aversion. In the following, we use $p\in\{1,\frac{1}{2},2\}$, corresponding to the log utility, a more risk-seeking, and a more risk-averse option, respectively. The graphs of these utility functions are plotted in Figure 2.
Figure 2: Utility functions $u_{p}$ for different powers $p$.
@@ 4.3.2 Path-dependent Penalties
In this section, we analyze our GAN-based approach for the penalized robust utility optimization problem of Equation 4 under the general path-dependent penalties of Equations 7 and 6. In contrast to the penalties of Equation 5 that we considered in Sections 4.1 and 4.2, these path-dependent penalties can be estimated well from observable quantities.
Moreover, observing the paths of trained generator and discriminator outputs $\pi^{\text{NN}}$ and $\mu^{\theta}$, $\Sigma^{\theta}$, in Figure 1 respectively, we notice the following: in the setting of Section 4.1.3 which included point-wise penalties as in Equation 5, the adversarial has learned to produce adversarial market parameters that almost stay constant, even in the fully robust case, where the adversarial was allowed to vary both drift and diffusion terms. This leads us to believe that the penalty (5) that was suggested in Guo et al. (2022) and which we have considered so far, was too strict. The penalty required the adversarial instantaneous drift and diffusion coefficients $\mu^{\theta}_{t}$ and $\Sigma^{\theta}_{t}$ to be close to $\tilde{\mu}$ and $\tilde{\Sigma}$, respectively, at all times during the training. Indeed, such penalties might not allow for much exploration.
Instead, in this section we consider path-wise penalties $R_{1},R_{2}$ that build on an observed data path, as introduced in Equations 7 and 6. In particular, we consider the quadratic covariation $\ell_{QCV}$ penalized via (6) (which is a constraint for the volatility) and the average relative return $l_{\text{ARR}}$ penalized via (7) (which is a constraint for the drift). Since we work with discretized processes on a time grid $0=t_{0}<\ldots<t_{N}=T$, we approximate the quadratic covariation and average relative return functionals $\ell_{QCV}$ and $l_{\text{ARR}}$, with $\hat{l}_{\text{QCV}}(S):=\sum_{n=1}^{N}(S_{t_{n}}-S_{t_{n-1}})^{2}$ and $\hat{l}_{\text{ARR}}(S):=\frac{1}{|B|}\sum_{b\in B}S_{T}(\omega_{b})/S_{0}(\omega_{b})$, respectively, where $B$ is a batch of paths generated by the adversarial market.
In synthetic settings, where we do not have observed data as reference, we assume a reference market (12) with constant parameters $\tilde{\mu},\tilde{\Sigma}$ and replace $S^{\mathrm{obs}}$ by $S^{\tilde{\mu},\tilde{\Sigma}}$ in Equations 6 and 7. In particular, this corresponds to a Black–Scholes model as reference market. The corresponding reference values are

### v2

@@ 4 Experiments
We start with experiments in easy settings, where an explicit solution exists in Section˜4.1. Then, in Sections˜4.2 and 4.3, we introduce friction, different utility functions and more realistic path-dependent penalization. In all of these cases, explicit solutions are not known. Therefore, in Section˜4.3.3, we introduce different metrics to evaluate the performance of our model. Moreover, in Section˜4.4, we leverage our GAN-based algorithm in an experimental study of small, proportional trading costs on optimal investments in robust and non-robust utility optimization.
@@
4.1 Friction-less Settings with Explicit Solutions
In the following examples we always use stock markets with risk free interest rate $r=1.5\%$, step size $\delta_{t}=\frac{1}{65}$ and $65$ steps. All stocks start at the initial value $S_{0}=1$ and are given by an Itô-diffusion driven by one or multiple Brownian motions. We sample a total of 200200K paths of Brownian motions, which are used for all experiments in this section such that there is no random influence due to different samples across different algorithm evaluations, with an $80/20$ train/test split.
We use the explicit solution for the market volatility $\Sigma^{*}$ (and $\mu^{*}$ if the drift is also robustified) to get an evaluation metric that computes the expected utility of either the learned or the explicit investment strategy against the explicit solution for the market. Let us define
@@ 4.1 Friction-less Settings with Explicit Solutions
then we compute $E_{u}(\pi,\Sigma^{*},\mu^{*})$ for the learned strategy $\pi=\pi^{\text{NN}}$ and for the explicit strategy $\pi=\pi^{*}$ (where $\mu^{*}=\mu$ in case the drift is fixed) and compare them via their relative error
@@ 4.1.1 Stock Market with 1 Risky Asset & Robust Volatility
$R_{\lambda_{1}}(\mathbb{P}_{\mu,\sigma})=\mathbb{E}_{\mathbb{P}_{\mu,\sigma}}\left[\lambda_{1}\int_{0}^{T}(\sigma_{t}-\tilde{\sigma})^{2}\,dt\right],$
where $\tilde{\sigma}\in\mathbb{R}_{+}$ is a constant reference value for the volatility (this could for example be estimated from historical data), and $\lambda_{1}$ is the scaling factor for the penalization. In this case there exists an explicit solution for the optimal actor and market strategy, which was given in Guo et al. (2022, Section 7.1). In particular, the optimal investment strategy and volatility are both constant and given as the solution of
@@
which we solve numerically to get a reference value. We use this setting to test and compare different models, since it offers an easy evaluation metric. In particular, we compute the expected utility (without penalty) when using the actor’s trained investment strategy and the explicit solution for the volatility (evaluation metric) and compare it to the expected utility when using the explicit solution for the actor and the market (reference evaluation metric). The reason for this type of evaluation procedure is that our ultimate goal is to achieve a good robust investment strategy from the GAN training, which should perform well in the harshest possible market environment.
@@ 4.1.2 Stock Market with Multiple Risky Assets & Robust Volatility
$R_{\lambda_{1}}(\mathbb{P}_{\mu,\sigma})=\mathbb{E}_{\mathbb{P}_{\mu,\sigma}}\left[\lambda_{1}\int_{0}^{T}\lVert\Sigma_{t}-\tilde{\Sigma}\rVert_{F}^{2}\,dt\right],$ (14)
where $\Sigma_{t}=\sigma_{t}\sigma_{t}^{\top}$. The utility function is $u(x)=\ln(x)$. To derive the explicit solution of the (continuous-time) optimization problem (4), we first use Itô’s lemma to derive the continuous dynamics of $\ln(X_{T}^{\pi,\mu,\sigma})$. After dropping the martingale terms, (4) simplifies to
@@
We follow (Guo et al., 2022, Section 7.2) and compute the instantaneous derivatives of the integrand with respect to $\pi_{t}$ and $\Sigma_{t}$ and set them to 0. This leads (after some rearrangement) to the equation systems
@@
The explicit solutions $\pi^{*},\Sigma^{*}$ can then be computed by first substituting $\Sigma$ in the second equation and numerically solving it for $\pi$ to obtain $\pi^{*}$ and then using this to compute $\Sigma^{*}$.
@@
      negatively correlated, asymmetric setting: assets’ volatilities are $0.15$ and $0.35$, their correlation is 0.9-0.9, leading to the reference volatility
@@
In all cases, our model achieves a very small relative error compared to the explicit strategy $\pi^{*}$ satisfying $\operatorname{err}_{rel}^{\Sigma^{\star},\mu}(\pi^{\text{NN}},\pi^{\star})<0.2\%$.
@@ 4.1.3 Robustifying Volatility and Drift
In this section, we robustify against uncertainty in the drift in addition to uncertainty in the volatility. Analogously to the robustification w.r.t. the volatility of Section˜4.1.2, we choose a reference value $\tilde{\mu}$ for the drift and consider the joint penalty function (5). Then, for log-utility $u=\ln$, we can compute the explicit solution as in Section 4.1.2. In particular, we first apply Itô’s lemma to (4), then drop the martingale terms and compute the instantaneous derivative of the integrand with respect to $\pi_{t},\Sigma_{t}$ and $\mu_{t}$. Setting them to 0 yields
$\Sigma=\frac{\pi\pi^{\top}}{4\lambda_{1}}+\tilde{\Sigma}\quad\text{and}\quad\Sigma\pi=(\mu-r)\quad\text{and}\quad\pi=2\lambda_{2}(\tilde{\mu}-\mu),$
which can be solved numerically for the unique explicit solution $\pi^{*},\Sigma^{*},\mu^{*}$. Therefore, we can again compute the relative errors $\operatorname{err}_{rel}^{\Sigma^{\star},\mu^{*}}(\pi^{\text{NN}},\pi^{\star})$.
@@
and use balanced penalty weights $\lambda_{1}=\lambda_{2}=1$ (cf. Equation˜5). In all combinations, we find a relative error $\operatorname{err}_{rel}^{\Sigma^{\star},\mu^{*}}(\pi^{\text{NN}},\pi^{\star})<0.2\%$.
Moreover, we get similar results when replacing the recurrent neural network (RNN) structures of generator and discriminator by standard feed-forward neural networks (FFNN) (that are reused for each time step). This was anticipated, since the optimal explicit strategy is constant and therefore not path dependent. In Figure 1, we show sample paths of the trained trading strategy evaluated against the explicit solution $\Sigma^{\star},\mu^{\star}$ for the market, and against the trained discriminator for the market on the test dataset. We see that in both cases, the trained strategy resembles the explicit strategy quite well. Interestingly, we further observe that the trained discriminator produces volatilities fairly similar to the explicit solution $\Sigma^{\star}$ (bottom left panels in Figure˜1. This is not the case for adversarially learned drift values that differ from $\mu^{\star}$ (top right panels in Figure˜1).
Figure 1: Test sample paths of the trained trading strategy evaluated against the explicit solution for the market (left) and against the trained discriminator for the market (right). A FFNN is used for generator and discriminator.
@@ 4.2 Settings with Proportional Transaction Costs
Since explicit solutions $\pi_{c}^{*},\Sigma_{c}^{*}$ (and $\mu^{*}$) for the problem including transaction costs are not known, we instead use the explicit solution for the corresponding problem without transaction costs (problem (15)), which can be considered as the best strategy that can be derived explicitly. In contrast to this limitation of closed-form solutions, a simple adjustment of the training framework allows us to train our GAN model in the market with transaction costs. As before, we compute the relative errors $\operatorname{err}_{rel}^{\Sigma^{\star},\mu^{*}}(\pi^{\text{NN}},\pi^{\star})$, where market friction is in place. Clearly, the explicit market parameters are, in general, not the optimizers of the problem with friction, however, since our goal is to achieve a robust strategy, it is still reasonable to test the strategies against these (maybe suboptimal) market parameters. We note again that negative values correspond to superior performance of the GAN-based method.
We consider the same settings for the reference volatility $\tilde{\Sigma}$ and reference drift $\tilde{\mu}$ as in Section˜4.1.3. For small transaction costs $c_{\text{prop}}=0.1\%$, we get $\operatorname{err}_{rel}^{\Sigma^{\star},\mu^{*}}(\pi^{\text{NN}},\pi^{\star})\in(-2.3\%,-0.3\%)$, which indicates slight outperformance of the (friction-aware) GAN model compared to the explicit solution $\pi^{\star}$ of the friction-less case. In particular, we see that taking friction into account improves the performance of the strategy, but that the difference is not very large due to the small size of the transaction costs. For larger transaction costs $c_{\text{prop}}=1\%$, this effect becomes considerably stronger, with relative errors resulting in $\operatorname{err}_{rel}^{\Sigma^{\star},\mu^{*}}(\pi^{\text{NN}},\pi^{\star})\in(-165.3\%,-10.5\%)$.
@@ 4.3 Realistic Settings for Generic Penalized Robust Utility Optimization
In this section, we analyze what we consider the most realistic setting for penalized robust utility optimization: We robustify against uncertainty in both the volatility and the drift, in a market with transaction costs, where we consider general power utilities corresponding to various preference structures including but not limited to log-preferences. Moreover, we study the path-wise penalties $R_{1}$ and $R_{2}$ of Equations˜7 and 6. These penalties are more realistic than the penalties of Sections˜4.1 and 4.2, since $R_{1}$ and $R_{2}$ can be estimated well from observed stock paths in contrast to the (empirically) unobservable instantaneous volatility and drift parameters that appear in Equation˜14.
In this general setting, no explicit solution to the robust utility optimization problem is known. Moreover, this setting differs in too many regards from the setting in Section˜4.1, where only log utility, no transaction costs and instantaneous penalization were considered, such that it is not reasonable to benchmark against the explicit solutions of Section˜4.1 anymore. Therefore, we introduce new metrics to evaluate model performance in Section˜4.3.3. Prior to that, we define general power utility functions and realistic path-wise penalties in Sections˜4.3.1 and 4.3.2.
@@ 4.3.1 Power Utility as Generalization of Log Utility
where $p\in\mathbb{R}_{\geq 0}\setminus\{1\}$. In the limit $p\to 1$ the $\log$-utility function is recovered, hence we write $u_{1}:=\ln$, and for $p=0$ $u_{p}$ equals the risk-neutral utility id1id-1. In particular, $p\in(0,1)$ leads to less risk aversion compared to the $\log$-utility function, while $p>1$ leads to more risk aversion. In the following, we use $p\in\{1,\frac{1}{2},2\}$, corresponding to the log utility, a more risk-seeking, and a more risk-averse option, respectively. The graphs of these utility functions are plotted in Figure˜2.
Figure 2: Utility functions $u_{p}$ for different powers pp.
@@ 4.3.2 Path-dependent Penalties
In this section, we analyze our GAN-based approach for the penalized robust utility optimization problem of Equation˜4 under the general path-dependent penalties of Equations˜7 and 6. In contrast to the penalties of Equation˜5 that we considered in Sections˜4.1 and 4.2, these path-dependent penalties can be estimated well from observable quantities.
Moreover, observing the paths of trained generator and discriminator outputs $\pi^{\text{NN}}$ and $\mu^{\theta}$, $\Sigma^{\theta}$, in Figure˜1 respectively, we notice the following: in the setting of Section˜4.1.3 which included point-wise penalties as in Equation˜5, the adversarial has learned to produce adversarial market parameters that almost stay constant, even in the fully robust case, where the adversarial was allowed to vary both drift and diffusion terms. This leads us to believe that the penalty (5) that was suggested in Guo et al. (2022) and which we have considered so far, was too strict. The penalty required the adversarial instantaneous drift and diffusion coefficients $\mu^{\theta}_{t}$ and $\Sigma^{\theta}_{t}$ to be close to $\tilde{\mu}$ and $\tilde{\Sigma}$, respectively, at all times during the training. Indeed, such penalties might not allow for much exploration.
Instead, in this section we consider path-wise penalties $R_{1},R_{2}$ that build on an observed data path, as introduced in Equations˜7 and 6. In particular, we consider the quadratic covariation $\ell_{QCV}$ penalized via (6) (which is a constraint for the volatility) and the average relative return $l_{\text{ARR}}$ penalized via (7) (which is a constraint for the drift). Since we work with discretized processes on a time grid $0=t_{0}<\ldots<t_{N}=T$, we approximate the quadratic covariation and average relative return functionals $\ell_{QCV}$ and $l_{\text{ARR}}$, with
$\hat{l}_{\text{QCV}}(S)_{i,j}$ $:=\sum_{n=1}^{N}(S_{t_{n}}^{i}-S_{t_{n-1}}^{i})(S_{t_{n}}^{j}-S_{t_{n-1}}^{j}),$ (18)
$\hat{l}_{\text{ARR}}(S)_{i}$ $:=\frac{1}{|B|}\sum_{b\in B}S_{T}^{i}(\omega_{b})/S_{0}^{i}(\omega_{b}),$ (19)
respectively, where $B$ is a batch of paths generated by the adversarial market.
In synthetic settings, where we do not have observed data as reference, we assume a reference market (12) with constant parameters $\tilde{\mu},\tilde{\Sigma}$ and replace $S^{\mathrm{obs}}$ by $S^{\tilde{\mu},\tilde{\Sigma}}$ in Equations˜6 and 7. In particular, this corresponds to a Black–Scholes model as reference market. The corresponding reference values are